Find an equation of the line that passes through the point (2, 3) and is perpendicular to the line
4x + 7y − 2 = 0

Answer by mananth(16949)

4 x + 7 y = 2
Find the slope of this line

7y=-4x +2
Divide by 7
y = - 4/ 7 x + 2/7
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = - 4/7

The slope of a line perpendicular to the above line will be the negative reciprocal 7/4
m1*m2=-1
The slope of the required line will be 1 3/4

m= 7/4 ,point ( 2 , 3 )
Find b by plugging the values of m & the point in
y=mx+b
3 = 7/ 2 + b
b= - 1/2
m= 7/4
The required equation is y = (7/ 4)x- 1/ 2
m.ananth@hotmail.ca